OpenStudy (anonymous):

What property is 5(2x)=(5•2)x Distributive, communitive or associative? I'll give a medal

OpenStudy (yanasidlinskiy):

Do you know what the difference between all of them? If you do..You should be fine:D

OpenStudy (zzr0ck3r):

if not use wiki.

OpenStudy (yanasidlinskiy):

Lol, zzr0ck3r. Wiki has an edit button;)

OpenStudy (anonymous):

is it distributive

OpenStudy (zzr0ck3r):

just google the definition.

OpenStudy (zzr0ck3r):

and no

OpenStudy (anonymous):

associative

OpenStudy (anonymous):

Communitive Property: The Commutative Property is the one that refers to moving stuff around. For addition, the rule is "a + b = b + a"; in numbers, this means 2 + 3 = 3 + 2. For multiplication, the rule is "ab = ba"; in numbers, this means 2×3 = 3×2. Associative Property: The Associative Property is the rule that refers to grouping. The rule is "a + (b + c) = (a + b) + c"; in numbers, this means 2 + (3 + 4) = (2 + 3) + 4. For multiplication, the rule is "a(bc) = (ab)c"; in numbers, this means 2(3×4) = (2×3)4. Distributive Property: The Distributive Property is easy to remember, if you recall that "multiplication distributes over addition". Formally, they write this property as "a(b + c) = ab + ac". In numbers, this means, that 2(3 + 4) = 2×3 + 2×4.

OpenStudy (anonymous):

so associative

OpenStudy (zzr0ck3r):

of ?

OpenStudy (anonymous):

multiplication

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I cant believe Im this old and still dont know the differences. ha. thanks a bunch

OpenStudy (anonymous):

because 5(2x) can have their parenthesis changed into (5*2)x and the answer will still be the same.